OpenStudy (anonymous):

The area of a square is equal to twice the perimeter of the square. Find the perimeter of the square. Choose one answer. a. 8 units b. 16 units c. 64 units d. 32 units

OpenStudy (anonymous):

Area = s^2 Perimeter = 4s Given area = 2*perimeter This implies, S^2 = 8s Cancelling s on both sides we get s=8 Therefore perimeter = 4*8 = 32 sq units

OpenStudy (anonymous):

thanks do you know this one Suppose a gas tank is in the shape of a cylinder. The tank is 10 feet long and 5 feet in diameter. Find the amount of metal used to construct the tank. Choose one answer. a. 128.5 ft2 b. 257 ft2 c. 392.6 ft2 d. 196.3 ft2

OpenStudy (anonymous):

i dont think i know how to do that equation

OpenStudy (anonymous):

its ok thanks anyways

OpenStudy (anonymous):

Np c:

OpenStudy (anonymous):

ok what about this one Five times the width of a certain rectangle exceeds twice its length by one inch, and three times its length is ten more than its perimeter. Find the width of the rectangle. Choose one answer. a. 21 inches b. 10 inches c. 42 inches d. 52 inches